# SKYSURF IX - The Cosmic Optical and Infrared Background from Integrated Galaxy Light Measurements

### v1

@@ 3.2 Star-Galaxy Separation
Without color information for all SKYSURF fields, only object sizes and single filter magnitudes can be used for star-galaxy classification. For star-galaxy separation we use the apparent magnitude and R50 parameters provided by ProFound. R50 is defined as the radius in arcseconds within which half of an object’s integrated flux is contained. Figure 7 shows the apparent magnitude versus plane for the ACSWFC F606W filter. The stars are shown as blue points and follow a nearly constant versus magnitude relationship below AB 20 mag. Between AB 20-22 mag, we fit a double Gaussian model in to the histogram of as shown in Figure 8 with the dividing boundaries marked. The stars form a narrow locus until AB 20 mag when their PSF begins to over-saturate the detector and their measured sizes increase. We use the width of the normal distribution occupied by the stars to define a gentle slope starting from AB 20.25-21.75 mag and ending at the mag boundary at which the populations blend together at faint magnitudes. While the minimum size between the two populations shown in Figure 8 occurs at -0.8, visual inspections of objects fainter than AB 21.25 mag confirm that objects between the displayed boundary and the galaxy locus confirmed that these objects are galaxies. Star-galaxy separation is done before performing galactic extinction corrections. At bright magnitudes we define a steeper slope from the size-magnitude relationship of the stars that forks upwards when the stars begin to over-saturate the detector. This is done independently for each filter. We also use the measured width of the stellar population, , to define a boundary between real objects and artifacts, which are denoted by green points in Figure 7.
@@ 4 Addressing HST’s Observing Biases
Finally, HST data is also prone to over-densities caused by excessive fragmentation of nearby galaxies during source detection and especially nearby low surface brightness objects, dwarf galaxies, globular clusters, and in some instances faulty readouts. While most of the area in such fields remains usable, these fields may contain unrealistic object densities at faint magnitudes. We thus elect to perform a third filtering step. We measure the number of objects brighter than AB 26 mag classified as galaxies in subregions (400x400 pixels for ACSWFC and WFC3UVIS, 100x100 for WFC3IR. The highest object density subregion for each field is stored. We removed fields containing subregions more dense than that 99th percentile of the total subregion density distribution. An example of a problematic field which we are able to identify this way is shown in Figure 13.
@@ 5 Galaxy Number Counts and the IGL
Here we discuss and present the SKYSURF galaxy number counts for the HST filter sets used in this analysis. In Figure 14 we show the galaxy number counts (left), integrated galaxy light (centre), and the normalized sum of the IGL (right) for the F606W filter as an example. To convert from galaxy number counts to a measurement of the EBL, we use the following standard equations
@@
The best spline fit to the data in Figure 14 is shown as a dashed black line alongside a subset of 100 Monte-Carlo spline fits to the data as described in Section 5.1. The splines shown in Figure 14 use 10 degrees of freedom. We report an extrapolated IGL value where we integrate between AB 100 mag, which is indistinguishable from integrating between . These values are reported in Table 6. We also report lower limits in Table 7 where we integrate to the faintest complete magnitude bin. When our galaxy number counts do not reach faint magnitudes, and/or the slope of the faint number counts is poorly defined, there can be significant differences between the extrapolated and non-extrapolated IGL. For example, in the F625W filter where the extrapolated IGL value of , and the lower limit ie . We thus elect to report a conservative mean value in with an uncertainty stretching between the on the respective bounds. These values are given in Tables 8 & 11.
@@ 5.1 Poisson Statistical Errors
For all spline fitting discussed in this section we use a weight on each data point where is the typical (where N is the number of galaxies in each half magnitude bin) Poisson statistical uncertainty on each data point. Integration under the spline fits to the central panel in Figure 14 allows us to obtain our IGL measurements. When deriving the Poisson error, the magnitude zero-point corrections provided in Tables 12 and 13 are used without any uncertainty in order to isolate the Poisson error. We independently shift each data point by its 1 uncertainty, then an error-weighted spline is fit to the perturbed data and the IGL value is obtained using the same integration process described above in Equations 1 & 2. This process is repeated 10,000 times where the median, 17th and 83rd percentiles of the resultant distribution are used to define a percentage uncertainty due to Poisson statistical error only.
@@ 5.2 Cosmic Variance Errors
Calculating a cosmic variance error for our HST number counts is non-trivial because the area changes with depth as seen in Table 5. As such, we calculate a fractional CV error for each SKYSURF galaxy count bin. This calculation is done using the cosvararea function in the celestial R software package555https://github.com/asgr/celestial/. The function takes a number of independent sight lines, area, and redshift range to calculate a volume and return the CV error as a percentage using the formulation from Driver & Robotham (2010). We assume each SKYSURF field is fully independent for the purposes of calculating the number of independent sight lines. We obtain the redshift range from the DEVILS redshift catalogs in the D03 and D10 fields. We take the 17th and 83rd percentile of the redshift distribution for objects from AB 18-24 mag as the inputs for cosvararea. This is done separately for each band in the ugrizYJH set given that the redshift distribution changes with wavelength. We obtain the CV uncertainty for both WAVES and DEVILS using the cosvararea function as well, where the value is constant regardless of magnitude given all magnitude bins cover the same area and number of independent sight lines.
We again perform Monte-Carlo analysis by shifting the number counts up/down according to their fractional cosmic variance errors again following a normal distribution with a mean of 1 and standard deviation of . As noted, the data from each survey (WAVES, DEVILS, and SKYSURF) are shifted systematically and independently for each data set. An error weighted spline is fit to the data and integrated. This process is repeated 10,000 times where the median, 17th and 83rd percentiles of the resultant distribution are used to define a percentage uncertainty due to CV error.
@@ 5.3 Magnitude Zero Point Corrections
Magnitude zero point errors will also affect the data systematically. We choose to color-correct number count data from WAVES and DEVILS onto the HST filter system which produces a systematic magnitude offset for both surveys. We calculate the color correction terms using the ProSpect filterTranBands function. As input, the function takes a filter, input filter, reference filter, and redshift range. The function returns the 3 terms used in the linear mappings. ProSpect comes with a long list of pre-calculated apparent magnitudes in each filter set which are used as inputs for the following equation
where , , and are terms returned by filterTranBands. is the apparent magnitude from the input filter, is the magnitude for a reference filter, and is the desired magnitude in the target filter. For example, when correcting WAVES Y-band data to HST’s WFC3IR F105W filter, the target, input, and reference filters are WFC3IR F105W, Y VISTA, and H VISTA, respectively. Using 1,000 of the ProSpect provided magnitudes, the final zero-point correction values and their uncertainties are provided in Tables 12 and 13.
Since we merge number counts from filters that exist in both ACSWFC and WFC3UVIS such as F606W, we take the mean of the two correction values and their respective uncertainties to correct ground based data to the HST filter system. We then take the correction and shift the WAVES and DEVILS number count magnitudes according to a normal distribution by uncertainties on the colour transformation parameters. We do this 10,000 times to obtain a spread of IGL values and use the 17th and 83rd percentiles to obtain an uncertainty, see Figure 15. The magnitude zero point corrections and their uncertainties are listed in the appendix.
@@ 5.4 Fitting Error
For all of the spline fitting in this work, we use a spline with 10 degrees of freedom by default. We choose to introduce a fitting error by fitting the data with splines of order 8-12 and taking the best fit IGL value for each order of spline. This is consistent with what has been done in previous works, see Driver et al. (2016b); Koushan et al. (2021). To stop the spline fits from diverging at the bright we ignore any galaxy number count bins with 50 or fewer galaxies. See Figure 22 (rows 3-4) where bright galaxy number count bins in WAVES-g band (green points) could introduce diverging spline fits. For the F140W (Figure 25, row 2) filter where no ground-based data is available we simply ignore the brightest two magnitude bins which otherwise cause the splines to diverge.
@@ 6.1 Comparison to Previous IGL Studies
Here we compare our IGL measurements to previous studies. To make a fair comparison, when comparing to previous IGL works we magnitude zero-point correct our HST number counts to the ground-based ugrizYJH filter system. Over the last decade studies of the IGL have consistently improved due in large part to the accumulation of galaxy number count data reducing cosmic variance uncertainties. Beginning with Driver et al. (2016b), the compilation of number counts from different existing surveys produced IGL measurements with uncertainties between . Cosmic variance contributed a dominant 10% while zero-point uncertainties remained steady at 4.5%. These errors arise from the smaller area coverage of the Galaxy and Mass Assembly Survey (GAMA) (Driver & Robotham, 2010; Liske et al., 2015) relative to WAVES and a significant number of data sets brought to a common magnitude zero-point. In Koushan et al. (2021), the introduction of DEVILS (Davies et al., 2021) increased area coverage at faint magnitudes and thus reduced the number of supplementary data sets required, which reduced cosmic variance and zero-point uncertainties to bring the IGL uncertainty down to 5%. Our extrapolated IGL uncertainties are compared to previous IGL studies in Figure 18 (bottom).
@@
Within the respective errors, we find our measurements to be consistent with those of Carter et al. 2025 (submitted), though our IGL values are found to be slightly higher. This can likely be attributed to ProFound’s well documented tendency to extract more flux from faint sources compared to SourceExtractor Robotham et al. (2018).
@@
The 1.8% error imposed by CV in WAVES in commination with comparable magnitude zero-point uncertainties introduced by using a suite of HST filters limits our improvements over Koushan et al. (2021). Our most precise measurements come from HST filters where we cover significant area and magnitude zero-point corrections are small such as the F814W, F125W, and F160W filters.
From Figure 19 we see that extrapolated IGL measurements are consistently above that of previous works while still within each other’s 1 uncertainties. Our higher IGL values could most likely be attributed to ProFound’s tendency to extract slightly more flux from faint objects than SourceExtractor (Bertin & Arnouts, 1996) as found in Robotham et al. (2018). The HST data used in Driver et al. (2016b); Koushan et al. (2021) was processed using SourceExtractor. In this work, all photometry for DEVILS, WAVES, and HST is provided by ProFound. Nevertheless all measurements are consistent within their quoted errors.
@@ 6.2 Comparison to VHE and Direct Measurements of the COB
Finally, we compare our IGL measurements to different measurements of the COB in Figure 19 (right). Our measurements are consistent with estimates of the COB provided by VHE studies (Desai et al., 2017; Fermi-LAT Collaboration et al., 2018; Gréaux et al., 2024). VHE studies provide a probe of the COB independent of the systematic uncertainties discussed in Section 5. Our extrapolated IGL value at is also consistent with the latest direct measurements from New Horizons presented in Postman et al. (2024). Values of , , and from our IGL measurements, the direct measurements of Postman et al. (2024), and VHE measurements of Gréaux et al. (2024) no longer present statistically significant disagreement. Therefore, refining measurements of the COB will provide a better understanding of the remaining foreground flux that constitutes the difference between measurements of the IGL and direct measurements. As the COB and IGL become equivalent, there is merit in further reducing uncertainties in IGL the to in order to serve as constraint on models.
With diffuse and/or unresolved sources of the COB ruled out, the IGL contains all of the information necessary to reconstruct the history of UV-NIR photon production since recombination. Thus the COB can be used as a powerful constraint on models of galaxy formation and evolution as measurements become more precise. Previous works such as (Koushan et al., 2021) demonstrated that the IGL could be used as a constraint on the CSFH. At the time, measurement precision meant that only the normalization of existing CSFH models could be tested. With more precise measurements the shape of the CSFH and models of more complex processes within galaxies can be evaluated using the COB as a constraint.
Finally, we display our measurement of the integrated COB () alongside many different historical measurements, Figure 20. To obtain our measurement, we fit a spline with 7 degrees of freedom to the IGL data from this work, Driver et al. (2016b), and Koushan et al. (2021) between . We then perform the same Monte-Carlo analysis as was done for the IGL measurements discussed in Section 5.1 and integrate 10,000 times. The median, 17th, and 83rd percentiles serve as our measurement of the COB and its 1 uncertainties. We obtain a COB value of , an uncertainty of 2%.

### v2

@@ 3.2 Star-Galaxy Separation
Without color information for all SKYSURF fields, only object sizes and single filter magnitudes can be used for star-galaxy classification. For star-galaxy separation we use the apparent magnitude and R50 parameters provided by ProFound. R50 is defined as the radius in arcseconds within which half of an object’s integrated flux is contained. Figure 7 shows the apparent magnitude versus plane for the WFC3IR F125W filter. The stars are shown as blue points and follow a nearly constant versus magnitude relationship below AB 20 mag. Between AB 20-22 mag, we fit a double Gaussian model in to the histogram of as shown in Figure 8 with the dividing boundaries marked. The stars form a narrow locus until AB 20 mag when their PSF begins to over-saturate the detector and their measured sizes increase. We use the width of the normal distribution occupied by the stars to define a gentle slope starting from AB 20.25-21.75 mag and ending at the mag boundary at which the populations blend together at faint magnitudes. While the minimum size between the two populations shown in Figure 8 occurs at -0.8, visual inspections of objects fainter than AB 21.25 mag confirm that objects between the displayed boundary and the galaxy locus confirmed that these objects are galaxies. Star-galaxy separation is done before performing galactic extinction corrections. At bright magnitudes we define a steeper slope from the size-magnitude relationship of the stars that forks upwards when the stars begin to over-saturate the detector. This is done independently for each filter. We also use the measured width of the stellar population, , to define a boundary between real objects and artifacts, which are denoted by green points in Figure 7.
@@ 4 Addressing HST’s Observing Biases
HST data is also prone to over-densities caused by excessive fragmentation of nearby galaxies during source detection and especially nearby low surface brightness objects, dwarf galaxies, globular clusters, and in some instances faulty readouts. While most of the area in such fields remains usable, these fields may contain unrealistic object densities at faint magnitudes. We thus elect to perform a third filtering step. We measure the number of objects brighter than AB 26 mag classified as galaxies in subregions (400x400 pixels for ACSWFC and WFC3UVIS, 100x100 for WFC3IR. The highest object density subregion for each field is stored. We removed fields containing subregions more dense than that 99th percentile of the total subregion density distribution. An example of a problematic field which we are able to identify this way is shown in Figure 13.
@@ 5 Galaxy Number Counts and the IGL
Here we discuss and present the SKYSURF galaxy number counts for the HST filter sets used in this analysis. In Figure 14 we show the galaxy number counts (top right), integrated galaxy light (top left), goodness of fit (bottom left), and the normalized sum of the IGL (bottom right) for the F125W filter as an example. To convert from galaxy number counts to a measurement of the EBL, we use the following standard equations. These equations are used with the AB magnitude system. This mathematical derivation of the IGL from galaxy number counts follows the same methodology as Driver et al. (2016b); Koushan et al. (2021).
@@
The best spline fit to the data in Figure 14 is shown as a dashed black line alongside a subset of 100 Monte-Carlo spline fits to the data as described in Section 5.1. The splines shown in Figure 14 use 10 degrees of freedom. The weight for each data point is set solely by the Poisson uncertainty, . These weights are maintained when testing different sources of uncertainty. In Figure 14 Poisson, cosmic variance, and magnitude zero point uncertainties are tested simultaneously. The implementation of each is described in the following subsections. The same figures are shown for all other bands in the appendix, see Figures 24, 25, 26, 27, 28, 29, & 30. We report an extrapolated IGL value where we integrate between AB 100 mag, which is indistinguishable from integrating between . These values are reported in Table 6. We also report lower limits in Table 7 where we integrate to AB 15 mag. We do not include WAVES galaxy number counts brighter than AB 15 mag in the fitting to force a Euclidean behaviour of the bright number counts. This is maintained throughout all IGL measurements. When our galaxy number counts do not reach faint magnitudes, and/or the slope of the faint number counts is poorly defined, there can be significant differences between the extrapolated and non-extrapolated IGL. For example, in the F625W filter where the extrapolated IGL value of , and the lower limit ie . We thus elect to report a conservative mean value in with an uncertainty stretching between the on the respective bounds. These values are given in Tables 8 & 11.
@@ 5.1 Poisson Statistical Errors
For all spline fitting we use a weight on each data point where is the typical (where N is the number of galaxies in each half magnitude bin) Poisson statistical uncertainty on each data point. To the Poisson error we also add an error based on our field selection criteria to account for the significant sampling bias in our HST data. For each galaxy magnitude bin we take the difference between the untrimmed and trimmed galaxy number counts shown in Figure 12 left & right, respectively. The square root of this value is added in quadrature to the Poisson error. The untrimmed HST galaxy number counts are shown in Figure 14, as downward facing red triangles, representing absolute upper limits. These are not used in fitting and illustrate the effectiveness of our field selection criteria. When deriving the Poisson error, the magnitude zero-point corrections provided in Tables 12 and 13 are used without any uncertainty in order to isolate the Poisson error. We independently shift each data point according to a normal distribution with a mean of 1 and standard deviation given by , the fractional Poisson statistical uncertainty. Then a spline is fit to the perturbed data and the IGL value is obtained using the same integration process described above in Equations 1 & 2. This process is repeated 10,000 times where the median, 17th and 83rd percentiles of the corresponding distribution are used to define a percentage uncertainty due to Poisson statistical error only.
@@ 5.2 Cosmic Variance Errors
Calculating a cosmic variance error for our HST number counts is non-trivial because the area changes with depth as seen in Table 5. As such, we calculate a fractional CV error for each SKYSURF galaxy count bin. This calculation is done using the cosvararea function in the celestial R software package555https://github.com/asgr/celestial/. The function takes a number of independent sight lines, area, and redshift range to calculate a volume and return the CV error as a percentage using the formulation from Driver and Robotham (2010). We assume each SKYSURF field is fully independent for the purposes of calculating the number of independent sight lines. We obtain the redshift range from the DEVILS redshift catalogs in the D03 and D10 fields. We take the 17th and 83rd percentile of the redshift distribution for objects from AB 18-24 mag as the inputs for cosvararea. This is done separately for each band in the ugrizYJH set given that the redshift distribution changes with wavelength. We obtain the CV uncertainty for both WAVES and DEVILS using the cosvararea function as well, where the value is constant regardless of magnitude given all magnitude bins cover the same area and number of independent sight lines.
Cosmic variance is a systematic uncertainty that affects each survey independently. We do not include the CV uncertainty in the weight assigned to each data point in the spline fit and maintain the weights as . To properly address CV as a systematic error, we shift each entire set of number counts up/down according to their fractional cosmic variance errors again following a normal distribution with a mean of 1 and standard deviation of . As SKYSURF number counts are collated based on exposure time, the CV uncertainty at the faint end is larger than at the bright end, which is reflected in the relatively large spread of splines at the faint end in Figure 14. This process is repeated 10,000 times where the median, 17th and 83rd percentiles of the resultant distribution are used to define a percentage uncertainty due to CV error.
. We comment on discontinuities and discrepancies between our measurements in Section 5.6.
@@ 5.3 Magnitude Zero Point Corrections
Magnitude zero point (ZPT) errors will also affect the data systematically. We ZPT correct number count data from WAVES and DEVILS onto the HST filter system which produces a systematic magnitude offset for both surveys. We calculate the color correction terms using the ProSpect filterTranBands function. As input, the function takes a filter, input filter, reference filter, and redshift range. ProSpect comes with a long list of pre-calculated apparent magnitudes in each filter set which are used as inputs for the following equation
where , , and are terms returned by filterTranBands. To determine the best solution for and , & , the quantity is minimised for all possible linear combinations of the reference and input facility filters, see Robotham (2020). is the apparent magnitude from the input filter, is the magnitude for a reference filter, and is the desired magnitude in the target filter. For example, when correcting WAVES Y-band data to HST’s WFC3IR F105W filter, the target, input, and reference filters are WFC3IR F105W, Y VISTA, and H VISTA, respectively. Using 1,000 of the ProSpect provided magnitudes, the final zero-point correction values and their uncertainties are provided in Tables 12 and 13.
Since we merge number counts from filters that exist in both ACSWFC and WFC3UVIS such as F606W, we take the mean of the two correction values and their respective uncertainties to correct ground based data to the HST filter system. To derive the magnitude zero point uncertainties, we shift the WAVES and DEVILS number count magnitudes according to a normal distribution with a mean given by the magnitude offset and the standard deviation as the magnitude offset error. We do this 10,000 times to obtain a spread of IGL values and use the 17th and 83rd percentiles to obtain an uncertainty, see Figure 15. The magnitude zero point corrections and their uncertainties are listed in the appendix.
@@ 5.4 Fitting Error
For all of the spline fitting in this work, we use a spline with 10 degrees of freedom by default. We choose to introduce a fitting error by fitting the data with splines of order 8-12 and taking the best fit IGL value for each order of spline. This is consistent with what has been done in previous works, see Driver et al. (2016b); Koushan et al. (2021). To stop the spline fits from diverging at the bright we ignore any galaxy number count bins with 50 or fewer galaxies. See Figure 25 where bright galaxy number count bins in WAVES-g band (green points) could introduce diverging spline fits. For the F140W (Figure 30, bottom) filter where no ground-based data is available we simply ignore the brightest two magnitude bins which otherwise cause the splines to diverge.
5.5 Dependencies on Ecliptic and Galactic Latitude
We also test our counts for selection biases that may arise based on ecliptic latitude or galactic latitude. HST’s standard processing steps and our own have performed sky subtractions, but we verify that when splitting our sample by ecliptic latitude, there are no significant changes. As Zodiacal light intensity depends strongly on ecliptic latitude, if sky subtraction imposes a systematic error it would be reflected in the IGL measurements when isolating SKYSURF fields in different ecliptic latitude bands. For both galactic and ecliptic latitude, we test the robustness of our sample by measuring the IGL in wide bands, respectively. As the small number of fields in different regions can introduce divergent integrations, we elect to use lower limits by integrating across the completeness limits of our data as was done in Table 7.
SKYSURF excluded all fields within of the galactic plane. We further test our IGL measurements by evaluating the IGL in wide galactic latitude bands from 10-30, 30-50, 50-70, and 70-90 degrees, respectively, Figure 17 The width of each galactic latitude selection is denoted by the dashed grey line through each point. During this process, we continue to include Poisson and Cosmic Variance, and magnitude zero-point uncertainties to account for the decreased number of fields. We find no significant dependencies on galactic latitude as shown in Figure 17. This indicates that galactic extinction corrections do not introduce a systematic error at lower galactic latitudes. The number density of stars also increases at low galactic latitude, and thus we can deduce that our star-galaxy separation does not introduce significant uncertainty to our IGL measurements.
SKYSURF did not make any selections based on ecliptic latitude, and thus we break our tests into wide intervals starting from zero. The intensity of the Zodical light varies significantly with ecliptic latitude and in principle could bias photometry. However, HST’s standard processing pipeline performs initial sky subtraction during calibration (Windhorst et al., 2022). We perform additional sky subtraction as part of our mosaic building and source detection. In Figure 18 we show IGL measurements in different wide bands. We find no consistent trend with ecliptic latitude and can conclude that HST’s and our own Zodical foreground subtraction is sufficient to provide consistent photometry.
@@ 6.1 Comparison to Previous IGL Studies
Here we compare our IGL measurements to previous studies. To make a fair comparison, when comparing to previous IGL works we magnitude zero-point correct our HST number counts to the ground-based ugrizYJH filter system. Over the last decade studies of the IGL have consistently improved due in large part to the accumulation of galaxy number count data reducing cosmic variance uncertainties. Beginning with Driver et al. (2016b), the compilation of number counts from different existing surveys produced IGL measurements with uncertainties between . Cosmic variance contributed a dominant 10% while zero-point uncertainties remained steady at 4.5%. These errors arise from the smaller area coverage of the Galaxy and Mass Assembly Survey (GAMA) (Driver and Robotham, 2010; Liske et al., 2015) relative to WAVES and a significant number of data sets brought to a common magnitude zero-point. In Koushan et al. (2021), the introduction of DEVILS (Davies et al., 2021) increased area coverage at faint magnitudes and thus reduced the number of supplementary data sets required, which reduced cosmic variance and zero-point uncertainties to bring the IGL uncertainty down to 5%. Our extrapolated IGL uncertainties are compared to previous IGL studies in Figure 20 (bottom).
@@
Within the respective errors, we find our measurements to be consistent with those of Carter et al. 2025 (submitted), though our IGL values are found to be slightly higher. This can be attributed to ProFound’s well documented tendency to extract more flux from faint sources compared to SourceExtractor Robotham et al. (2018). In particular, the flux extraction difference between the two tools is primarily a result of the segment dilation approach employed by ProFound. The mechanics of the dilation approach are discussed extensively in Robotham et al. (2018). Both extended low surface brightness and faint sources require more dilation steps until the flux extraction converges.
@@
The 1.2% error imposed by CV in WAVES in commination with comparable magnitude zero-point uncertainties introduced by using a suite of HST filters limits our improvements over Koushan et al. (2021). Our most precise measurements come from HST filters where we cover significant area and magnitude zero-point corrections are small such as the F814W, F125W, and F160W filters.
From Figure 21 we see that extrapolated IGL measurements are consistently above that of previous works while still within each other’s 1 uncertainties. Our higher IGL values could most likely be attributed to ProFound’s tendency to extract slightly more flux from faint objects than SourceExtractor (Bertin and Arnouts, 1996) as found in Robotham et al. (2018). The HST data used in Driver et al. (2016b); Koushan et al. (2021) was processed using SourceExtractor. In this work, all photometry for DEVILS, WAVES, and HST is provided by ProFound. Nevertheless all measurements are consistent within their quoted errors.
@@ 6.2 Comparison to VHE and Direct Measurements of the COB
Finally, we compare our IGL measurements to different measurements of the COB in Figure 21 (right). Our measurements are consistent with estimates of the COB provided by VHE studies (Desai et al., 2017; Fermi-LAT Collaboration et al., 2018; Gréaux_2024). VHE studies provide a probe of the COB independent of the systematic uncertainties discussed in Section 5. Our extrapolated IGL value at is also consistent with the latest direct measurements from New Horizons presented in Postman et al. (2024). Values of , , and from our IGL measurements, the direct measurements of Postman et al. (2024), and VHE measurements of Gréaux_2024 no longer present statistically significant disagreement. Therefore, refining measurements of the COB will provide a better understanding of the remaining foreground flux that constitutes the difference between measurements of the IGL and direct measurements. As the COB and IGL become equivalent, there is merit in further reducing uncertainties in IGL the to in order to serve as constraint on models.
With significant diffuse and/or unresolved sources of the COB ruled out, the IGL contains all of the information necessary to reconstruct the history of UV-NIR photon production since recombination. Thus the COB can be used as a powerful constraint on models of galaxy formation and evolution as measurements become more precise. Previous works such as (Koushan et al., 2021) demonstrated that the IGL could be used as a constraint on the CSFH. At the time, measurement precision meant that only the normalization of existing CSFH models could be tested. With more precise measurements the shape of the CSFH and models of more complex processes within galaxies can be evaluated using the COB as a constraint.
Finally, we display our measurement of the integrated COB () alongside many different historical measurements, Figure 22. To obtain our measurement, we fit a spline with 10 degrees of freedom to the extrapolated IGL data from this work, Driver et al. (2016b), and Koushan et al. (2021) between . We then perform the same Monte-Carlo analysis as was done for the IGL measurements discussed in Section 5.1 and integrate 10,000 times. The median, 17th, and 83rd percentiles serve as our measurement of the COB and its 1 uncertainties. We obtain a COB value of , an uncertainty of 2%.